OpenStudy (kkutie7):

Now suppose this student told you that she had sissolved the unknown in 250.00 mL, but only used 10.00 mL of that 250.00 mL solution to do the precipitation with 8-HQ. How many grams of Al were in the ORIGINAL 250.00 mL solution (i.e., in the original unknown sample)?

OpenStudy (kkutie7):

[\[Al^{3+}+3C_{9}H_{6}(OH)N:\rightarrow Al[C_{9}H_{6}(O^{-})N:]_{3}\downarrow + 3H^{+}\]

OpenStudy (australopithecus):

Not sure what it means by 8-HQ is that what the actual question says? what you need to realize is that concentration is consistent, for example if you take a portion of a solution that portion will have the same concentration as its parent solution.

OpenStudy (kkutie7):

8-HQ is the same as \[3C_{9}H_{6}(OH)N:\]

OpenStudy (kkutie7):

I also have this information: Aluminum precipitated= .0238grams of Al and Al(HQ)3 complex precipitated= 1.2176 g

OpenStudy (australopithecus):

You didnt provide enough information to solve this problem, you need a concentration or a quantity of 8-HQ that you used to cause complete precipitation of the unknown compound in the 10mL solution of it to figure out the unknowns concentraiton

OpenStudy (kkutie7):

I don't have that

OpenStudy (australopithecus):

Then this question cannot be answered

OpenStudy (australopithecus):

The experiment has not been completed thus there is no data to allow you to determine your answer

OpenStudy (kkutie7):

It has to be. its on my prelab. this is all I have: Calculate the formula weight of the Al [C9H6ON]3 complex that will be formed in this experiment.=459.385. A student performed a gravimetric analysis of an unknown solid sample by precipitating the aluminum using 8-HQ, forming the Al(HQ)3 complex as you will do in this lab. If the precipitate, after subtracting the crucible and microfilter, weighed 1.2176 g, what is the mass (in g) of aluminum in the precipitate?= .0238grams of Al

OpenStudy (kkutie7):

@Australopithecus

OpenStudy (australopithecus):

Ok you do have enough information to answer this

OpenStudy (kkutie7):

phew! good. show me how to go about this please

OpenStudy (australopithecus):

Is 0.0238 your guess as to how much aluminium is in your precipitate?

OpenStudy (kkutie7):

thats is what I calculated yes

OpenStudy (australopithecus):

So you know how much product you produced: 1.2176 g Using, Gravimetric formula \[Grams\; of\; compound*\frac{Molecular\; Mass\; of\; Element}{Molecular\; Mass\; Of\; Compound} = Grams\; of\; Element\] If this is what you did you are fine

OpenStudy (kkutie7):

no I did this: \[\frac{1.2176g of Al[C_{9}H_{6}(O^{-})N:]_{3}}{1}*\frac{1mole of Al[C_{9}H_{6}(O^{-})N:]_{3}}{459.3852g of Al[C_{9}H_{6}(O^{-})N:]_{3}}*\frac{1mole Al}{3moles of Al[C_{9}H_{6}(O^{-})N:]_{3}} \]

OpenStudy (australopithecus):

0.0238g*(250.0mL/10.0mL) = Grams of Al in original solution Make up a ratio in this case 250mL/10mL = the amount you need to scale up Notice that the number gets larger which is what you expected.

OpenStudy (kkutie7):

\[\frac{1.2176g of Al[C_{9}H_{6}(O^{-})N:]_{3}}{1}*\frac{1mole of Al[C_{9}H_{6}(O^{-})N:]_{3}}{459.3852g of Al[C_{9}H_{6}(O^{-})N:]_{3}}\] \[*\frac{1mole Al}{3moles of Al[C_{9}H_{6}(O^{-})N:]_{3}}\]

OpenStudy (australopithecus):

Whenever a solution is diluted or a portion of it taken from the stock solution you just need to create a ratio to figure out the amount in the original solution

OpenStudy (kkutie7):

ok I get that 250is divided by 10

OpenStudy (australopithecus):

yeah, makes sense right?

OpenStudy (australopithecus):

Let me check your calculation

OpenStudy (kkutie7):

ok

OpenStudy (kkutie7):

so you are saying that the answer would be .595 grams of Al?

OpenStudy (australopithecus):

http://www.webqc.org/molecular-weight-of-Al(C9H6ON)3.html so, 459.4317g/mol for the compound Al has a molecular mass of 26.98g/mol so using gravimetric formula: (26.98g/mol/459.4317g/mol)*1.2176g = 0.0715g http://www.wolframalpha.com/input/?i=%2826.98g%2Fmol%2F459.4317g%2Fmol%29*1.2176g+%3D+

OpenStudy (kkutie7):

gravimetric formula?

OpenStudy (kkutie7):

why do you multiply by 1.2176? whats wrong with the way I did it?

OpenStudy (kkutie7):

also need help with this I know its weight of solute/ weight of solution*100 Supposed another student told you he calculated the number of grams of Al in the original 250.00 mL solution to be 1.788 g, and that he prepared that solution by dissolving 4.7400 g of a solid unknown in water. What is the wt % Al3+ in the original solid sample? which is which?

OpenStudy (australopithecus):

I have no idea what you did, mine just uses the ratio of Al to compound, that percentage can be used to figure out how many grams of compound are made up by that specific element. here is to show that other people use this formula https://ca.answers.yahoo.com/question/index?qid=20100202072330AAJTPqC I wish I could help you but I have to eat and spend time with friends etc. I will be able to help you later tonight maybe if I can. I recommend opening a new question though. I hope I was helpful